# Homeomorphisms of continua through projective Fraïssé limits

### v1

@@ 1. $(\hexstar)_{W}$
We argue that the moreover clause follows from the main part of the lemma. Pick $m_{+}\in M$, $m^{\prime}_{+}\in M^{\prime}$ such that $f(m_{+})=f^{\prime}(m^{\prime}_{+})$ is an endpoint of $L$. Let $\alpha:N\to M$ be an epimorphism which maps the endpoint $n_{-}$ ($n_{+}$, resp.) of $N$ to $m_{-}$ ($m_{+}$, resp.) and similarly, $\alpha^{\prime}:N^{\prime}\to M^{\prime}$ is such that $n_{-}^{\prime}$ goes to $m_{-}^{\prime}$, and $n_{+}^{\prime}$ is sent to $m_{+}^{\prime}$. Since
$f\circ\alpha(n_{+})=f(m_{+})=f^{\prime}(m^{\prime}_{+})=f\circ\alpha(n^{\prime}_{+})\textrm{ is an endpoint of }L,$
@@
(B.6) ${\rm tp}^{n_{+},N}(f\circ\alpha)={\rm tp}^{n^{\prime}_{+},N^{\prime}}(f^{\prime}\circ\alpha^{\prime}).$
On the other hand, $\alpha(n_{-})=m_{-}$ is an endpoint, so necessarily
${\rm tp}^{n_{-},N}(f\circ\alpha)={\rm tp}^{m_{-},M}(f),$
@@
${\rm tp}^{n^{\prime}_{-},N^{\prime}}(f^{\prime}\circ\alpha)={\rm tp}^{m^{\prime}_{-},M^{\prime}}(f^{\prime}),$
@@
(B.7) ${\rm tp}^{n_{-},N}(f\circ\alpha)={\rm tp}^{n^{\prime}_{-},N^{\prime}}(f^{\prime}\circ\alpha),$
Using (B.6), (B.7) we can apply the lemma to $f\circ\alpha$ and $f^{\prime}\circ\alpha^{\prime}$. Replacing the resulting $g$ and $g^{\prime}$ by $\alpha\circ g$ and $\alpha^{\prime}\circ g^{\prime}$, respectively, we are done. ∎

### v2

@@ 1. $({\rm R}2)$
Suppose that $I,J,J^{\prime}$, $f,f^{\prime}$ and $a_{J}\in J$, $a_{J^{\prime}}\in J^{\prime}$ are as in our assumptions. Pick $c_{J}\in J$, $c_{J^{\prime}}\in J^{\prime}$ such that $f(c_{J})=f^{\prime}(c_{J^{\prime}})$ is an endpoint of $I$. Let $\alpha:K\to J$ be an epimorphism which maps the endpoint $a_{K}$ ($b_{K}$, resp.) of $K$ to $a_{J}$ ($c_{J}$, resp.) and similarly, $\alpha^{\prime}:K^{\prime}\to J^{\prime}$ is such that $a_{K^{\prime}}$ goes to $a_{J^{\prime}}$, and $b_{K^{\prime}}$ is sent to $c_{J^{\prime}}$. Since
$f\circ\alpha(b_{K})=f(c_{J})=f^{\prime}(c_{J^{\prime}})=f\circ\alpha(b_{K^{\prime}})\textrm{ is an endpoint of }I,$
@@
(B.10) ${\rm tp}^{b_{K},K}(f\circ\alpha)={\rm tp}^{b_{K^{\prime}},K^{\prime}}(f^{\prime}\circ\alpha^{\prime}).$
On the other hand, $\alpha(a_{K})=a_{J}$ is an endpoint, so it is not hard to see that
${\rm tp}^{a_{K},K}(f\circ\alpha)={\rm tp}^{a_{J},J}(f),$
@@
${\rm tp}^{a_{K^{\prime}},K^{\prime}}(f^{\prime}\circ\alpha^{\prime})={\rm tp}^{a_{J^{\prime}},J^{\prime}}(f^{\prime}),$
@@
(B.11) ${\rm tp}^{a_{K},K}(f\circ\alpha)={\rm tp}^{a_{K^{\prime}},K^{\prime}}(f^{\prime}\circ\alpha^{\prime}),$
So by (B.10), (B.11) Lemma B.2 applies. Replacing the resulting $g$ and $g^{\prime}$ by $\alpha\circ g$ and $\alpha^{\prime}\circ g^{\prime}$, respectively, we are done. ∎